OpenStudy (anonymous):

What is the Mass of a hubcap that was thrown off a car at 25m/s2 with a Force of 50N ?

OpenStudy (amistre64):

doesnt force equate to mass times acceleration?

OpenStudy (anonymous):

Yes F=M *A

OpenStudy (amistre64):

then A = F/m

OpenStudy (anonymous):

They wanna find the mass not the force though

OpenStudy (amistre64):

LOL, M = f/a

OpenStudy (amistre64):

accursed cap lock button, how i loathe thee

OpenStudy (anonymous):

I just divide?

OpenStudy (amistre64):

well yeah, just make sure your units of measurement are the same not sure what a newton is in terms of m and s^2

OpenStudy (amistre64):

N = kg m/s^2 i believe which cancels out to kg

OpenStudy (anonymous):

I divided and got 0.5

OpenStudy (anonymous):

So 0.5 Kg?

OpenStudy (amistre64):

50/25 is not 1/2

OpenStudy (anonymous):

2

OpenStudy (amistre64):

ugh, this site is starting to go wonky on my screen

OpenStudy (anonymous):

Yu said M= f/a

OpenStudy (anonymous):

Not M = a/f

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

prove it